If y = 1, then (x+5)•y=x+5. Which property supports this statement?
Answer by jim_thompson5910(35256)

(x+5)•y=x+5

(x+5)•1=x+5

x+5 = x+5

We're using the multiplicative identity property to go from step 2 to step 3